how do you write john is 11 more than twice as old as bob the sum of their ages is 65. find the ages of each. as a equation
Answer by mananth(16946)

Bob = x
John = 2x+11
...
x+2x+11=65
3x+11=65
-11 -11
3x=54
/3
x=18 Bob's age
John = 2x +11= 2*18+11=47
...
47+18=65
...
m.ananth@hotmail.ca